A portable animal cage that weighs 9,856 N rests on the floor. How much work (in J) is required to lift it 7.9 m vertically, at constant speed?
1
Expert's answer
2020-03-02T09:06:37-0500

Applying Newton's second Law to calculate the magnitude of the vertical force.


"F_{Net}=m\;a=F-W"


Where.

  • The acceleration a
  • The mass of the cage m
  • The weight of the cage W=9,856
  • Applied force F


Remember that the acceleration is zero because it moves at constant velocity.


"a=0\;\frac{m}{s^{2}}"


Thus.


"m\;a=F-W \\ m\;0=F-W \\ F=W"


The external force applied to raise the cage to constant velocity


"F=9,856\;N"


To calculate the work we use the following equation.


"W=F\;d\;cos\theta"


Where.

  • The force applied is "F=9,856\;N"
  • Displacement is "d=7.9\;m"
  • The angle between displacement and force is "\theta=0^{0}"


Evaluating numerically.

"W=F\;d\;cos\theta \\W=9,856\;N\times 7.9\;m\times cos(0^{0}) \\W=77,862.4\;J"


The required work is "W=77,862.4\;J"